OpenStudy (anonymous):

A particle moving in a straight line with constant retardation of 3ms^-2 has an initial velocity of 10ms^-1. find after what time it returns to its starting point

OpenStudy (anonymous):

@Hero @dpaInc

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@SmoothMath @phi @radar

OpenStudy (anonymous):

time is 10/3?

OpenStudy (anonymous):

i dont understand ><

OpenStudy (anonymous):

can u solve it using s= 1/2at^2 + ut?

OpenStudy (anonymous):

what is S0?

OpenStudy (anonymous):

we dont know initial postition?

OpenStudy (anonymous):

so S = S0

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

returns to its starting point implies that displacement = 0 ?!

OpenStudy (anonymous):

yes 0 = ut+ 1/2at^2 solve for t is it?

OpenStudy (callisto):

I'm thinking in this way. Not sure if it is correct though. I did bad in Physics :(

OpenStudy (anonymous):

i got it :) thanks

OpenStudy (callisto):

welcome :)

OpenStudy (phi):

we could think like this: the particle goes right at v= 10 m/s , but slowing by -3 m/s^2 the time it takes to slow to zero is given by \[ \Delta v= a t\] or 10= 3t, t= 10/3 seconds we could argue symmetry to say the return will also take 10/3 seconds for a total time of 30/3 or 6 2/3 seconds or we could say, after 10/3 of a second the particle has gone s= 0.5a t^2 + vt meters s= 0.5(-3)*100/9 + 10*10/3 = 50/3 meters the return trip is the same distance, with the initial velocity zero: 50/3 = 0.5(-3)* t^2 --> t=10/3 or we could graph the problem,